OpenStudy (anonymous):

help me !!!!!!!!!!!!!!!!!!! MEDALLL Find f(-2) ,f(-0.5), and f (3) for f(x)+3/8x - 3

OpenStudy (aum):

f(x)+3/8x - 3 Is the + supposed to be = ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

typo

OpenStudy (aum):

\[ f(x) = \frac{3}{8x-3} ~~~~ \text{ or } ~~~~ f(x) = \frac{3}{8x} - 3 ? \]

OpenStudy (anonymous):

the second one

OpenStudy (aum):

\[ f(x) = \frac{3}{8x} - 3 \\ f(-2) = \frac{3}{8(-2)} - 3 = \frac{3}{-16} - 3 = \frac{-3}{16} - 3 = \frac{-3-48}{16} = -\frac{51}{16}\\ f(0.5) = ? \]